OpenStudy (anonymous):

Solve x2 + 4x – 12 = 0 by completing the square. Show your work for full credit.

myininaya (myininaya):

first step add 12 on both sides \[x^2+4x=12\] now we want to complete the square on the left hand side so we need to add (b/2)^2 (4/2)^2=2^2=4 we add that on one side we have to do the same to other side \[x^2+4x+4=12+4\] \[(x+2)^2=16\] take square root of both sides \[x+2=\pm 4\] now im pretty sure you can subtract 2 on both sides

OpenStudy (anonymous):

x^2 + 4x = 12 (x + 2)^2 -4 = 12 (x+2^2 =16 (x+2) = =4 or -4 x = x = 2 or -6